# Uniform Feasibility For Smoothed Backup Control Barrier Functions

### v1

@@ III-B Compact Sets
Finally, the strict positiveness of the Lie derivative implies $\nabla\tilde{h}_{\theta}(x)\neq 0$ on $\partial\tilde{\mathcal{S}}_{\theta}$, and thus 0 is a regular value of $\tilde{h}_{\theta}$. Then $\tilde{\mathcal{S}}_{\theta}$ is strictly safe based on Aubin’s strict invariance condition [aubin1991viability]; and by Theorem  1, $\tilde{h}_{\theta}$ is a CBF for all $\theta>\theta^{*}$. ∎
@@ III-C General Case
In this part we study a set of sufficient conditions implying similar results without relying on the compactness property. In particular, since $T_{\varepsilon}$ may be unbounded, the condition (17) is no longer sufficient to satisfy (20) as the uniform bounds (24)-(26) are not guaranteed. Therefore, we distinguish between the portion where the state remains within a bounded region and the portion where $\|x\|$ grows unbounded. We refer to the former as the core and the latter as the tail. This distinction is useful because the core behaves as a compact set, so the arguments and uniform bounds from the previous section already apply, whereas the tail requires additional growth conditions to ensure a uniform lower bound on $L_{F}\tilde{h}_{\theta}$.
To formalize this, we decompose $T_{\varepsilon}$ using a radius $R>0$:
$T_{\varepsilon}=(T_{\varepsilon}\cap B_{R})\cup(T_{\varepsilon}\cap B_{R}^{\rm c}),\quad B_{R}\triangleq\{x\in R^{n}~|~\|x\|\leq R\},$ (41)
where $B_{R}^{\rm c}$ denotes the complement of $B_{R}$. The set $T_{\varepsilon}\cap B_{R}$ forms the compact core, whereas $T_{\varepsilon}\cap B_{R}^{\rm c}$ forms the tail, which is unbounded for any $R>0$ (if $\hat{\mathcal{S}}$ is unbounded). Therefore, we introduce additional assumptions to limit the contribution of inactive terms in $L_{F}\tilde{h}_{\theta}$ for the tail.
@@
Assumption 1 essentially assures that the active part of $L_{F}\tilde{h}_{\theta}$ (the first term in (23)) overcomes the inactive part (the second term) on the tail part ($\forall x\in T_{\varepsilon}\cap B_{R}^{\rm c}$). In particular, Assumption 1 (a) implies that $w_{\theta,j}\to 0$ for all inactive constraints $j\in\mathcal{P}(x)$ as $\|x\|\to 0$ with an exponential rate. Therefore, it will dominate the polynomial growth of $L_{F}h_{j}$ in this portion (per Assumption 1 (b)). As a result, the inactive part in (23) goes to 0 as $\|x\|\to\infty$ or $\theta\to\infty$. At any fixed $x$, we can find a finite $\theta^{*}$ such that the inactive term in (23) falls within the strictly positive margins of the uniform bound of the active part (that is, $r_{\infty}$ as assumed by Assumption 1 (c)).
@@
Assumption 1 (a) holds if (i) there is a unique minimizer for $\hat{h}$ at the tail end; i.e., $\exists i^{*}\in\mathbb{I}_{N},\mathcal{A}(x)\equiv i^{*},\forall x\in T_{\varepsilon}\cap B_{R}^{\rm c}$, and (ii) the gap $\delta_{j}$ is radially coercive for inactive constraints; i.e., $\delta_{j}(x)\to\infty$ as $\|x\|\to\infty$ holds for all $j\in\mathcal{P}(x)$. Under these conditions the function
$\eta(r)\triangleq\min_{j\neq i^{*}}\min_{\|x\|=r}\Big(h_{j}(x)-h_{i^{*}}(x)\Big)$ (45)
is well-defined, and it satisfies requirements in Assumption 1 (a). Failing to satisfy either of these may result arbitrarily small gap $\delta_{j}\to 0$ as $\|x\|\to\infty$, which prevents the dominance of the active part in $L_{F}\tilde{h}_{\theta}$.
@@
where
$\theta^{*}_{\mathrm{tail}}\triangleq\frac{1}{\eta(R)}\log\!\Bigg(\frac{(N-1)\big(r_{\infty}+C(1+R)^{p}\big)}{r_{\infty}}\Bigg).$ (50)
@@
The strict positiveness of the Lie derivative implies $\nabla\tilde{h}_{\theta}(x)\neq 0$ on $\partial\tilde{\mathcal{S}}_{\theta}$, and thus 0 is a regular value of $\tilde{h}_{\theta}$. Then $\tilde{\mathcal{S}}_{\theta}$ is strictly safe based on Aubin’s strict invariance condition [aubin1991viability]; and by Theorem 2, $\tilde{h}_{\theta}$ is an eCBF for all $\theta>\theta^{*}$.
IV Connection to the Backup CBF Framework
We now connect the general results of Section III to the backup CBF setting introduced in Section II-C. Let a (open-loop) system $f\in C^{1}$ as in (1), a set $\mathcal{S}$ and a backup set $\mathcal{S}_{\rm b}$ defined with functions $h,h_{\rm b}\in C^{1}$, respectively. Suppose that a backup controller $k_{\rm b}\in C^{1}$ yields a closed-loop vector field $F_{\rm b}(x)=f(x,k_{\rm b}(x))\in C^{1}$, where the flow is denoted as $\phi(x_{0},\tau)$ with an initial condition $x(0)=x_{0}\in\mathcal{S}$. Recall from Section II-C
@@ IV Connection to the Backup CBF Framework
(i) Compactness.  Define the map $\Phi_{\tau}(x)=\phi(x,\tau)$. As the solution flow uniquely exists on $\tau\in[-T,T]$ for all $x\in\mathbb{R}^{n}$, $\Phi_{\tau}$ is injective, because no two different initial conditions can lead to a same point in its image (if $\Phi_{\tau}(x_{1})=\Phi_{\tau}(x_{2})$, then $x_{1}=x_{2}$). Thus, $\Phi_{\tau}:\mathbb{R}^{n}\mapsto\Phi_{\tau}(\mathbb{R}^{n})$ is bijective for all $|\tau|\leq T$, where $\Phi_{\tau}(\mathbb{R}^{n})$ is its image. Furthermore, since $F_{\rm b}\in C^{1}$, $\Phi_{\tau}$ is $C^{1}$ for all $|\tau|\leq T$. Thus, $\Phi_{\tau}$ is a $C^{1}$ diffeomorphism with a $C^{1}$ inverse $\Phi_{-\tau}$ for all $|\tau|\leq T$ [lee2003smooth]. This implies that $\mathcal{B}_{N}$ is the continuous image of the compact set $\mathcal{S}_{\rm b}$ under the inverse: $\mathcal{B}_{N}=\Phi_{-T}(\mathcal{S}_{\rm b})$, and therefore $\mathcal{B}_{N}$ is compact.
(ii) Safety. We use the semigroup property of the flow similar to [gurriet2020scalable]—but for strict safety. Take any $x\in\mathcal{B}_{N}$, then $\Phi_{T}(x)\in\mathcal{S}_{\rm b}$ by definition of $\mathcal{B}_{N}$. Since the backup system is strictly safe w.r.t. $\mathcal{S}_{\rm b}$, for every $s>0$ we have $\Phi_{T+s}(x)\in{\rm Int}(\mathcal{S}_{\rm b})$. Thus, for every $s>0$, $b_{N}(\Phi_{s}(x))=h_{\rm b}(\Phi_{T+s}(x))>0$, so $\Phi_{s}(x)\in{\rm Int}(\mathcal{B}_{N})$. This is exactly the safety definition in Definition 3. ∎
@@
Assumption 2 translates as the strict safety wherever $\partial\mathcal{S}$ becomes relevant for the backup trajectory. Together with the strict safety of $\mathcal{S}_{\rm b}$, this yields the strict safety of $\hat{\mathcal{B}}$, which is required to invoke Theorem 3. Note that the condition is required only on $\mathcal{C}$ because this set consists of the boundary points of the target safe set $\mathcal{S}$ and the backward-expanded set of $\mathcal{B}_{N}$, i.e., exactly the states at which the backup-based construction is applied. Any other boundary point of $\mathcal{S}$ that is outside $\mathcal{C}$ is not relevant for the formulation as $x\notin\mathcal{B}_{N}$ implies $\phi(x,T)\notin\mathcal{S}_{\rm b}$.
@@
Under MFCQ, Aubin’s strict invariance condition [aubin1991viability] reduces to the strict positiveness of the Lie derivative on boundary, cf. (17). We verify this inequality at an arbitrary $x\in\partial\hat{\mathcal{B}}$ for the same two cases as in the previous part: $i\in\mathcal{A}(x)$ for a $i<N$, or $N\in\mathcal{A}(x)$. Case $i<N$, as discussed in the previous part, Assumption 2 applies, and $L_{F_{\rm b}}b_{i}(x)>0$ for all $x\in\partial\hat{\mathcal{B}}$ when $i$-th constraint is active with $i<N$. Case $i=N$, means $\phi(x,T)\in\partial\mathcal{S}_{\rm b}$ and, since the backup closed loop is strictly safe w.r.t. $\mathcal{S}_{\rm b}$ and 0 is a regular value for $h_{\rm b}$, we have $L_{F_{\rm b}}h_{\rm b}(\phi(x,T))>0$. Thus (17) holds, and strict safety follows from Aubin’s strict invariance condition [aubin1991viability].
@@
Corollary 1 relies on the compactness of the backup set $\mathcal{S}_{\rm b}$; and we can still certify $\tilde{b}_{\theta}$ as a CBF even in the case $h$ defines an unbounded $\mathcal{S}$ (provided that Assumption 2 holds). This relaxes the stronger (global) compactness assumptions found in the literature; e.g., [gurriet2020scalable]. If the backup set $\mathcal{S}_{\rm b}$ is not compact, then we can still use Theorem 4 to certify $h$ as an eCBF, which would imply the feasibility of the safety filter (5) all the same. We leave this as a topic for future work.
V Conclusion

### v2

@@ III-B Compact Sets
Finally, the strict positiveness of the Lie derivative implies $\nabla\tilde{h}_{\theta}(x)\neq 0$ on $\partial\tilde{\mathcal{S}}_{\theta}$, and thus 0 is a regular value of $\tilde{h}_{\theta}$. Then $\tilde{\mathcal{S}}_{\theta}$ is strictly safe based on Aubin’s strict invariance condition [2]; and by Theorem  1, $\tilde{h}_{\theta}$ is a CBF $\forall\theta>\theta^{*}$. ∎
While Theorem 1 certifies $\tilde{h}_{\theta}$ as a CBF and thereby guarantees the existence of a class-$\mathcal{K}$ function $\alpha$ for which the safety constraint in (5) is feasible, it does not by itself prescribe such an $\alpha$. By leveraging the converse safety theorem, however, one can derive a condition that yields a lower bound for such a class-$\mathcal{K}$ function. In particular, a linear class-$\mathcal{K}$ function $\alpha(r)=\gamma r$ leads to a feasible safety constraint whenever $\gamma$ satisfies
$\gamma\geq\gamma^{*}\triangleq\frac{L_{\rm max}}{\varepsilon},~L_{\rm max}\triangleq\displaystyle\sup_{x\in\tilde{S}_{\theta}}\Big[\max\bigl(0,-L_{F}\tilde{h}_{\theta}(x)\bigr)\Big]$ (40)
with $\varepsilon$ as in Lemma 1, and $L_{\rm max}$ representing the worst rate of change of $\dot{\tilde{h}}_{\theta}$ on $\tilde{\mathcal{S}}_{\theta}$. Inside the tube $T_{\varepsilon}$, the proof of Theorem 3 already ensures $L_{F}\tilde{h}_{\theta}(x)>0$, so the CBF condition holds for any $\gamma>0$. Outside the tube, the bound (40) ensures that $\gamma\,\tilde{h}_{\theta}(x)\geq-L_{F}\tilde{h}_{\theta}(x)$.
Remark 4 (Computability of $\theta^{*}$ and $\gamma^{*}$).
All quantities entering the threshold $\theta^{*}$ in (33) and the bound (40) on $\gamma$ can be determined offline from the system data: the constants $\varepsilon$, $M$, $r$, and $d$ in $\theta^{*}$ follow from Lemma 1, and $L_{\rm max}$ can be evaluated numerically by sampling once $\theta$ is fixed. Together, these provide a recipe for a priori feasibility of the safety constraint in (5) uniformly over $\tilde{\mathcal{S}}_{\theta}$.
@@ III-C General Case
In this part we study a set of sufficient conditions implying similar results without relying on the compactness property. Since $T_{\varepsilon}$ may be unbounded, the uniform bounds (23)-(25) need not hold and (16) alone is not sufficient for (19). We therefore decompose $T_{\varepsilon}=(T_{\varepsilon}\cap B_{R})\cup(T_{\varepsilon}\cap B_{R}^{\rm c})$ with $B_{R}\triangleq\{x~|~\|x\|\leq R\}$ and $B_{R}^{\rm c}$ denotes the complement of $B_{R}$. The compact core (the first term in the set union) inherits the bounds of Section III-B, and we impose additional conditions on the unbounded tail end (the second term) to control inactive terms in $L_{F}\tilde{h}_{\theta}$.
@@
Assumption 1 ensures that, on the tail end, the active part of $L_{F}\tilde{h}_{\theta}$ (the first term in (22)) dominates the inactive part (the second term): (a) forces $w_{\theta,j}\to 0$ exponentially for inactive constraints, which overcomes the polynomial growth allowed by (b). As a result, the inactive part in (22) goes to 0 as $\|x\|\to\infty$ or $\theta\to\infty$. At any fixed $x$, we can find a finite $\theta^{*}$ such that the inactive term in (22) falls within the strictly positive margins of the uniform bound of the active part (that is, $r_{\infty}$ as assumed by Assumption 1 (c)).
@@
Assumption 1 (a) holds, for example, when there is a unique minimizer $i^{*}$ of $\hat{h}$ at the tail end and the inactive gaps are radially coercive ($\delta_{j}(x)\to\infty$ as $\|x\|\to\infty$); then $\eta(r)\triangleq\min_{j\neq i^{*}}\min_{\|x\|=r}\left(h_{j}(x)-h_{i^{*}}(x)\right)$ satisfies the requirement.
@@
where $\theta^{*}_{\mathrm{tail}}\triangleq\frac{1}{\eta(R)}\log\!\Bigg(\frac{(N-1)\big(r_{\infty}+C(1+R)^{p}\big)}{r_{\infty}}\Bigg)$.
@@
The strict positiveness of the Lie derivative implies $\nabla\tilde{h}_{\theta}(x)\neq 0$ on $\partial\tilde{\mathcal{S}}_{\theta}$, and thus 0 is a regular value of $\tilde{h}_{\theta}$. Then $\tilde{\mathcal{S}}_{\theta}$ is strictly safe based on Aubin’s strict invariance condition [2]; and by Theorem 2, $\tilde{h}_{\theta}$ is an eCBF $\forall\theta>\theta^{*}$.
IV Connection to the BCBF Framework
We now connect the general results of Section III to the BCBF setting introduced in Section II-C. Let a (open-loop) system $f\in C^{1}$ as in (1), a set $\mathcal{S}$ and a backup set $\mathcal{S}_{\rm b}$ defined with functions $h,h_{\rm b}\in C^{1}$, respectively. Suppose that a backup controller $k_{\rm b}\in C^{1}$ yields a closed-loop vector field $F_{\rm b}(x)=f(x,k_{\rm b}(x))\in C^{1}$, where the flow is denoted as $\phi(x_{0},\tau)$ with an initial condition $x(0)=x_{0}\in\mathcal{S}$. Recall from Section II-C
@@ IV Connection to the BCBF Framework
(i) Compactness.  Define the map $\Phi_{\tau}(x)=\phi(x,\tau)$. As the solution flow uniquely exists on $\tau\in[-T,T]$ for all $x\in\mathbb{R}^{n}$, $\Phi_{\tau}$ is injective, because no two different initial conditions can lead to a same point in its image (if $\Phi_{\tau}(x_{1})=\Phi_{\tau}(x_{2})$, then $x_{1}=x_{2}$). Thus, $\Phi_{\tau}:\mathbb{R}^{n}\mapsto\Phi_{\tau}(\mathbb{R}^{n})$ is bijective for all $|\tau|\leq T$, where $\Phi_{\tau}(\mathbb{R}^{n})$ is its image. Furthermore, since $F_{\rm b}\in C^{1}$, $\Phi_{\tau}$ is $C^{1}$ for all $|\tau|\leq T$. Thus, $\Phi_{\tau}$ is a $C^{1}$ diffeomorphism with a $C^{1}$ inverse $\Phi_{-\tau}$ for all $|\tau|\leq T$ [13]. This implies that $\mathcal{B}_{N}$ is the continuous image of the compact set $\mathcal{S}_{\rm b}$ under the inverse: $\mathcal{B}_{N}=\Phi_{-T}(\mathcal{S}_{\rm b})$, and therefore $\mathcal{B}_{N}$ is compact.
(ii) Safety. We use the semigroup property of the flow similar to [9]—but for strict safety. Take any $x\in\mathcal{B}_{N}$, then $\Phi_{T}(x)\in\mathcal{S}_{\rm b}$ by definition of $\mathcal{B}_{N}$. Since the backup system is strictly safe w.r.t. $\mathcal{S}_{\rm b}$, for every $s>0$ we have $\Phi_{T+s}(x)\in{\rm Int}(\mathcal{S}_{\rm b})$. Thus, for every $s>0$, $b_{N}(\Phi_{s}(x))=h_{\rm b}(\Phi_{T+s}(x))>0$, so $\Phi_{s}(x)\in{\rm Int}(\mathcal{B}_{N})$. This is exactly the safety definition in Definition 3. ∎
@@
Assumption 2 translates as the strict safety wherever $\partial\mathcal{S}$ becomes relevant for the backup trajectory. Together with the strict safety of $\mathcal{S}_{\rm b}$, this yields the strict safety of $\hat{\mathcal{B}}$, which is required to invoke Theorem 3. Note that the condition is required only on $\mathcal{C}$ because this set consists of the boundary points of the target safe set $\mathcal{S}$ and the backward-expanded set of $\mathcal{B}_{N}$, i.e., exactly the states at which the backup-based construction is applied. Any other boundary point of $\mathcal{S}$ that is outside $\mathcal{C}$ is not relevant for the formulation as $x\notin\mathcal{B}_{N}$ implies $\phi(x,T)\notin\mathcal{S}_{\rm b}$. We note that Assumption 2, together with the other hypotheses of Corollary 1, can be verified directly from the system data, either analytically or numerically; the example in Section V illustrates such a verification on a grid.
@@
Under MFCQ, Aubin’s strict invariance condition [2] reduces to the strict positiveness of the Lie derivative on boundary, cf. (16). We verify this inequality at an arbitrary $x\in\partial\hat{\mathcal{B}}$ for the same two cases as in the previous part: $i\in\mathcal{A}(x)$ for a $i<N$, or $N\in\mathcal{A}(x)$. Case $i<N$, as discussed in the previous part, Assumption 2 applies, and $L_{F_{\rm b}}b_{i}(x)>0$ for all $x\in\partial\hat{\mathcal{B}}$ when $i$-th constraint is active with $i<N$. Case $i=N$, means $\phi(x,T)\in\partial\mathcal{S}_{\rm b}$ and, since the backup closed loop is strictly safe w.r.t. $\mathcal{S}_{\rm b}$ and 0 is a regular value for $h_{\rm b}$, we have $L_{F_{\rm b}}h_{\rm b}(\phi(x,T))>0$. Thus (16) holds, and strict safety follows from Aubin’s strict invariance condition [2].
@@
We note that conditions required by Corollary 1 can be verified directly from the system data, either analytically or numerically; the example in Section V illustrates such a verification on a grid. Beyond certifying $\tilde{b}_{\theta}$ as a CBF, Corollary 1 brings the practical recipe of Section III-B into the BCBF setting: once $\theta>\theta^{*}$ is fixed, a lower bound $\gamma^{*}$ can be computed from (40) for $\alpha(r)=\gamma r$. In the next section, we illustrate this procedure on a numerical example.
@@
Corollary 1 relies on the compactness of the backup set $\mathcal{S}_{\rm b}$; and we can still certify $\tilde{b}_{\theta}$ as a CBF even in the case $h$ defines an unbounded $\mathcal{S}$ (provided that Assumption 2 holds). This relaxes the stronger (global) compactness assumptions found in the literature; e.g., [9]. If the backup set $\mathcal{S}_{\rm b}$ is not compact, then we can still use Theorem 4 to certify $h$ as an eCBF, which would imply the feasibility of the safety constraint in (5) all the same. We leave this as a topic for future work.
V Example
Figure 1: (Top) Sets $\mathcal{S}$, $\mathcal{S}_{\rm b}$, $\hat{\mathcal{B}}$ and $\tilde{\mathcal{B}}_{\theta}$ for the example, and simulation trajectories with safety filter (5) with $\hat{b}$ (brown) and $\tilde{b}_{\theta}$ (orange) for $\alpha(r)=\gamma r$. (Middle) Feasibility of the safety constraint over time for $\hat{b}$ with $\gamma\in\{0.01,0.1,0.25,0.5,1\}$: all $\gamma<1$ exhibit intervals of infeasibility. (Bottom) Feasibility of the safety constraint over time for $\tilde{b}_{\theta}$: every run is feasible, consistent with the a priori certificate $\gamma^{*}>0$.
We illustrate the method on a double integrator
$\dot{x}_{1}=x_{2},\quad\dot{x}_{2}=u,\quad|u|\leq u_{\max}=2,$ (56)
and the safe set
$\mathcal{S}=\{x\in\mathbb{R}^{2}~|~h(x)\geq 0\},\quad h(x)=x_{1}+2x_{2}+1.$ (57)
Note that $\mathcal{S}$ is unbounded, and $h$ is not a CBF for (56) due to the input bound. This motivates the BCBF.
Backup controller and backup set. We use a linear backup controller $k_{\rm b}(x)=-x_{1}-2x_{2}$, and to obtain a $C^{1}$ backup closed-loop, we saturate $k_{\rm b}$ smoothly:
$F_{\rm b}(x)=\begin{bmatrix}x_{2}\\ u_{\max}\tanh\!\bigl((-x_{1}-2x_{2})/u_{\max}\bigr)\end{bmatrix}$. (58)
Solving the Lyapunov equation with $F_{\rm b}$ linearized around the origin and Q=IQ=-I yields $P=\begin{bmatrix}3/2&1/2\\ 1/2&1/2\end{bmatrix}$. We define the (compact) backup set as the Lyapunov sublevel set
$\mathcal{S}_{\rm b}=\{x\in\mathbb{R}^{2}~|~h_{\rm b}(x)\geq 0\},\quad h_{\rm b}(x)=\rho-x^{\top}Px,$ (59)
with $\rho=0.1$. This choice guarantees $\mathcal{S}_{\rm b}\subset\mathcal{S}$, $|k_{\rm b}(x)|<u_{\max}$ on $\mathcal{S}_{\rm b}$, and ensures that $F_{\rm b}$ is strictly safe w.r.t. $\mathcal{S}_{\rm b}$.
BCBF construction. With horizon $T=2$ and discretization step $\Delta\tau=0.1$, we construct the functions $\{b_{i}\}_{i=1}^{N}$ as in (9) with $N=21$, and define $\hat{b}$, $\hat{\mathcal{B}}$ as in (49). The safety filter (5) with the nonsmooth $\hat{b}$ enforces the $N$ simultaneous constraints (10). Existing nonsmooth BCBF formulations do not provide a constructive procedure for selecting the class-$\mathcal{K}$ functions $\alpha_{i}$, leaving the user to tune them by trial and error until a configuration yielding a feasible safety constraint is found. To illustrate the consequences of this ad hoc selection, we simulate the safety filter ($k_{\rm des}=-u_{\rm max}$) with $\alpha(r)=\gamma r$ for $\gamma\in\{0.01,0.1,0.25,0.5,1\}$ from the initial condition $x_{0}=(-2.5,3)$; the feasibility of safety constraints in each run is reported over time in the middle panel of Fig. 1. All choices with $\gamma<1$ yield intervals of infeasibility (the trajectory for $\gamma=1$ is shown in the top panel).
Smoothed BCBF. We replace the pointwise minimum with the soft-min $\tilde{b}_{\theta}$ as in (51). Verifying the hypotheses of Corollary 1: $\mathcal{S}_{\rm b}$ is compact, 0 is a regular value of $h_{\rm b}$, $F_{\rm b}\in C^{1}$, $F_{\rm b}$ is strictly safe w.r.t. $\mathcal{S}_{\rm b}$. Assumption 2 is checked numerically on a grid: at boundary points $x\in\partial\hat{\mathcal{B}}$ where an active constraint $i<N$ corresponds to $\phi(x,\tau_{i})\in\mathcal{C}$, we evaluate $L_{F_{\rm b}}b_{i}(x)$ and find it strictly positive throughout, confirming $L_{F_{\rm b}}h(y)>0$ on the sampled portion of $\mathcal{C}$. The case $i=N$ follows directly from the strict safety of $\mathcal{S}_{\rm b}$, which yields $L_{F_{\rm b}}h_{\rm b}(\phi(x,T))>0$ at such points.
We set $\varepsilon=\rho$ so that $T_{\varepsilon}=\hat{\mathcal{B}}$. The quantities $M,r,d$ from Lemma 1 and $L_{\rm max}$ from (40) are evaluated numerically on a uniform grid over a box containing $\hat{\mathcal{B}}$. Trajectories and sensitivities needed for $b_{i}$ and gradients $\nabla b_{i}$ are obtained by forward-integrating $F_{\rm b}$ together with the variational equation $\dot{S}=(\partial F_{\rm b}/\partial x)\,S$, $S(0)=I$, using a fixed-step Runge–Kutta scheme [^4] Related code are in https://github.com/anilalan-umich/smooth-bcbf. This yields $M\approx 2.01$, $r\approx 0.027$, and $d\approx 0.005$, which via (33) give $\theta^{*}\approx 1451$; we set $\theta=1524$, which gives a very close approximation of $\hat{\mathcal{B}}$ with $\tilde{\mathcal{B}}_{\theta}$. The resulting $L_{\rm max}$ is numerically zero as the backup controller can enforce a strict increase of $\tilde{b}_{\theta}$ everywhere on $\tilde{\mathcal{B}}_{\theta}$. This means that (40) is satisfied with any $\gamma>0$. Crucially, this certifies that any linear class-$\mathcal{K}$ function renders the safety constraint in (5) feasible on $\tilde{\mathcal{B}}_{\theta}$, a conclusion that cannot be provided for the classical nonsmooth BCBF formulation a priori. This is demonstrated in the bottom panel of Fig. 1, where the safety filter (5) with $\tilde{b}_{\theta}$ is simulated from the same initial condition $x_{0}$ for each $\gamma\in\{0.01,0.1,0.25,0.5,1\}$ and the constraint remains feasible throughout in every run; the closed-loop trajectory for $\gamma=1$ is shown in the top panel.
VI Conclusion